how do I solve these formulas? First I am to plug in the given values and then solve remaining variable.
A = 1/2 bh, b = 23 and h = 91
and
F = ma for F = 920. 1 and a = 136
Please explain because I am lost...
Thank you

Answer by Alan3354(69443)

how do I solve these formulas? First I am to plug in the given values and then solve remaining variable.
A = 1/2 bh, b = 23 and h = 91
and
F = ma for F = 920. 1 and a = 136
======================================
1st, formulas are not solved. Equations are solved.
----------------------
A = 1/2 bh, b = 23 and h = 91
Area = (1/2)bh = (1/2)*23*91
Area = 2093/2
Area = 1046.5
--------------------
and
F = ma for F = 920. 1 and a = 136
----
Sub the values given for the variables.
920 = m*136
Divide both sides by 136
m = 920/136
m = 115/17
m =~ 6.7647